Today's big event was the joining of the inner and outer frames and the mounting of our cryostat mass replacement, the so-called cryodummy. Basically, because of all the work the gondola crew did on mounting electronics and other hardware when the two were apart, now that everything is back together they are ready to hang and start doing scan tests!
On the receiver side, we had a partially botched fridge cycle (our 1K fridge didn't cycle at all), so that limited the amount of data we could take. Still, we managed to get some. On the downside, one of our DfMUX boards let out some of its magic smoke (from a capacitor). After a little investigation, the cause was found and it will be fixed forthwith. The fridges shall be cycled again tonight (with the issue that prevented last night's cycle found and corrected), so we should be good to go again tomorrow.